# stanhf: HistFactory models in the probabilistic programming language Stan

### v1

@@ I Introduction
Interpreting the results of a physics experiment requires a statistical model. This model describes the relationship between a theory and potential observations, including statistical and systematic errors that could affect the outcome. In collider physics, experiments are often based on counting the numbers of events in bins of a histogram. We present a new way to build and analyze these statistical models, based on the probabilistic programming language Stan [1, 2]. These computer representations of a statistical model preserve and communicate an experimental analysis and enable it to be modified or recasted.
To understand the form of these models, we start from the Poisson likelihood for data binned in a histogram or several histograms
@@
where and represent individual multiplicative and systematic effects for bin contribution . Lastly, the parameters might be constrained by prior knowledge or auxiliary experiments. Thus, combining these changes to Eq. 1, we arrive at
@@
Previously, the HistFactory specification was developed to represent these models [3]. This is a declarative specification of a statistical model of the form Eq. 5 — the modifiers and auxiliary term in Eq. 5 can be built from a set of predeclared choices. The specification was written in xml language and was intended to be interpreted by the RooFit/RooStats framework based on ROOT histogram functionality [4, 5, 6]. More recently, a json schema was designed to express HistFactory models in json [7, 8].
@@
We present a complementary tool to the successful pyhf and HistFactory based on Stan [1, 2]. Stan is an imperative probabilistic programming language with automatic differentiation [18], enabling models to be analyzed in a Bayesian framework using Hamiltonian Monte Carlo (HMC; see e.g. refs. [19, 20]). This is an efficient algorithm for obtaining a sample-representation of the posterior distribution. Stan started development in 2011, and is used in industry [21, 22] and academia, including cognitive, social and political sciences, medical statistics, and astrophysics [23, 24, 25].
Because Stan is an imperative language, it is more flexible and expressive than declarative specifications. This enables richer and more detailed modeling, as the modeling choices defined in the HistFactory declarative specification can be tweaked and adapted in the Stan language. This is a major difference from pyhf and bayesian_pyhf. Although Stan was designed primarily for Bayesian modeling and computation, frequentist statistics, such as confidence intervals and p-values, are available in stanhf through an interface with pyhf.
@@ II Quickstart
The process is summarized in Fig. 1 for the HistFactory model examples/model.json.
@@ III.1 Structure of Stan model
Stan is a strongly typed and imperative probabilistic programming language. Stan programs, including those created by stanhf such as that in Fig. 1, define a statistical model through special blocks.
@@
We declare the data in the data block; here, data includes experimental data, as well as nominal event numbers, and hyperparameters that define modifiers and constraint terms. We do not define data here; it is written to a separate json file. In the context of high-energy physics, the natural separation between a model and observed data in Stan makes data-blinding straight-forward [28].
@@
Unknown parameters, associated with modifiers, are declared in the parameters block. These parameters are marginalzed or profiled in an analysis.
@@
Finally, we simulate data from the mode in the generated quantities block. This simulated data can be used to compute posterior predictive distributions.
@@ III.2 Choices of prior
Stan is agnostic about whether constraint terms are part of the prior or likelihood, since it is only their product that matter. Thus, we can interpret stanhf models as having implicit flat priors for parameters between any bounds placed on them in the HistFactory model, and constraint terms that are part of the likelihood.
@@ III.6 Validation and testing
Checking that the parameter names and sizes agree; see stanhf.Converter.validate_param_names.
@@
The log-likelihods are evaluated at randomly chosen points found by perturbing the suggested initial choices by noise. See stanhf.Converter.validate_target.
@@ IV Bayesian inference
The stanhf CLI builds an executable Stan program and writes associated data files. We briefly describe how to use them for Bayesian inference. For concreteness, we look here at the examples/model.json model shown in Fig. 1. In this model, a signal is characterized by a strength parameter and corresponds to the background-only model. First, we must build the model and sample from it,
@@
We can plot posterior distributions. For example, to show the one-dimensional posterior pdf {mdframed}[ backgroundcolor=white, linewidth=1pt,linecolor=gray, innerleftmargin=-1pt, innertopmargin=-3pt, innerbottommargin=-6pt, innerrightmargin=4pt]
Lastly, the Bayes factor [37] and Bayes factor surface [38] can be computed using a Savage-Dickey ratio [39]. The Bayes factor surface as a function of relative to the background-only model could be found by, {mdframed}[ backgroundcolor=white, linewidth=1pt,linecolor=gray, innerleftmargin=-1pt, innertopmargin=-3pt, innerbottommargin=-6pt, innerrightmargin=4pt]
![[Uncaptioned image]](x2.png)
@@ V Frequentist inference
As well as using stanhf for Bayesian inference using the powerful HMC functionality of Stan, we can harness the expressive power of Stan models in a frequentist setting. stanhf provides classes and functions to use a Stan model and optimizer in place of a pyhf model or native pyhf optimizer in the statistical inference package of pyhy.infer, which uses results in ref. [40].
As an example, we again use examples/model.json, which declares that mu is the parameter of interest (POI). Using pyhf, we can plot the statistic as a function of : {mdframed}[ backgroundcolor=white, linewidth=1pt,linecolor=gray, innerleftmargin=-1pt, innertopmargin=-3pt, innerbottommargin=-6pt, innerrightmargin=4pt]
![[Uncaptioned image]](x3.png)
@@
VI Conclusions
We introduced stanhf — a new way to analyze and develop statistical models based on histogrammed data. The code transpiles HistFactory models into Stan, a probabilistic programming language. This allows state-of-the-art, scalable optimiztion and sampling using automatic differentiation and HMC. Because Stan is an imperative language, HistFactory modeling choices can be tweaked and adapted, as desired.
Stan is part of an ecosystem of Bayesian analysis software, allowing MC chains to be scrutinized for convergence and plotted, and even allowing models to be reran and analyzed online. Lastly, to ensure that the frequentist toolbox remains available, we provide an interface between Stan models and the inference machinery of pyhf.

### v2

@@ I Introduction
Interpreting the results of a physics experiment requires a statistical model. This model describes the relationship between a theory and potential observations, including statistical and systematic errors that could affect the outcome. In collider physics, experiments are often based on counting the numbers of events in bins of a histogram. We present a new way to build and analyze these statistical models, based on the probabilistic programming language (PPL) Stan [1, 2]. These computer representations of a statistical model preserve and communicate an experimental analysis and enable it to be modified or recasted.
More generally, PPLs are languages that are designed for statistical modeling (see e.g., refs. [3, 4]). Like most languages, they support random number generation and their outputs can be probabilistic; significantly, however, they are designed to make it easy to express statistical models and perform statistical inference. Besides Stan, examples of PPLs include WinBUGS [5], Turing.jl [6], PyMC [7], Edward [8, 9], pyro/numpyro [10, 11], WinBUGS [5] and RooFit/RooStats [12, 13, 14]. In the context of high-energy physics, PPLs were recently explored in refs. [15, 16, 17, 18, 19].
To understand the form of statistical models for these counting experiments, we start from the Poisson likelihood for data binned in a histogram or several histograms
@@
where indexes individual systematic effects, and and represent individual multiplicative and systematic effects for bin , contribution and effect . Lastly, the parameters might be constrained by prior knowledge or auxiliary experiments. Thus, combining these changes to Eq. 1, we arrive at
@@
On the other hand, rather than binning events into histograms, one could model the events using a probability density function (pdf). Suppose that we observed events with a discriminating variable, e.g. invariant mass, . Rather than binning with respect to the discriminating variable as in Eq. 1, we write
where the pdf models the discriminating variable for each event. Each observed event is marked by its value of the discriminating variable, and thus Eq. 6 is known as a marked Poisson model.
Previously, the HistFactory specification was developed to represent binned models [20]. This is a declarative specification of a statistical model of the form Eq. 5 — the modifiers and auxiliary term in Eq. 5 can be built from a set of predeclared choices. The specification was written in xml language and was intended to be interpreted by the RooFit/RooStats framework based on ROOT histogram functionality [12, 13, 14]. More recently, a json schema was designed to express HistFactory models in json [21, 22].
@@
We present a complementary tool to the successful pyhf and HistFactory based on Stan [1, 2]. Stan is an imperative probabilistic programming language with automatic differentiation [31], enabling models to be analyzed in a Bayesian framework using Hamiltonian Monte Carlo (HMC; see e.g. refs. [32, 33]). This is an efficient algorithm for obtaining a sample-representation of the posterior distribution. Stan started development in 2011, and is used in industry [34, 35] and academia, including cognitive, social and political sciences, medical statistics, and astrophysics [36, 37, 38]. As we build on HistFactory, we focus on binned models, though unbinned models such as Eq. 6 could be programmed in Stan.
Because Stan is an imperative language, it is more flexible and expressive than declarative specifications. This enables richer and more detailed modeling, as the modeling choices defined in the HistFactory declarative specification can be tweaked and adapted in the Stan language. This is a major difference from pyhf and bayesian_pyhf. Although Stan was designed primarily for Bayesian modeling and computation, frequentist statistics, such as confidence intervals and p-values, are available in stanhf through an interface with pyhf. Stan aims to offer state-of-the-art performance and supports computation on GPUs [39]. Performance comparisons to JAX [40], a possible computational backend in pyhf, indicate that Stan outperforms JAX on the CPU but that JAX outperforms Stan on the GPU [41].
@@ II Quickstart
The process is summarized for simplified example HistFactory models in Figs. 1 and 2. The simplified models include a normalization factor (denoted by normfactor in the HistFactory specification) and a normalization systematic uncertainty (denoted by normsys in the HistFactory specification), respectively. They can be found in the stanhf source code at examples/normfactor.json and examples/normsys.json, respectively.
@@ III.1 Structure of Stan model
Stan is a strongly typed and imperative probabilistic programming language. Stan programs, including those created by stanhf, define a statistical model through special blocks. We present example Stan programs produced from the simplified examples examples/normfactor.json and examples/normsys.json in Figs. 1 and 2, respectively.
@@
We declare the data in the data block; here, data includes experimental data, as well as nominal event numbers, and hyperparameters that define modifiers and constraint terms. We do not define data here; it is written to a separate json file. In the context of high-energy physics, the natural separation between a model and observed data in Stan makes data-blinding straight-forward [44], as observed data can be substituted without altering the model or model file.
@@
Unknown parameters, associated with modifiers, are declared in the parameters block. These parameters are marginalized or profiled in an analysis.
@@
Finally, we simulate data from the model in the generated quantities block. This simulated data can be used to compute posterior predictive distributions.
@@ III.2 Choices of prior
Stan is agnostic about whether auxiliary terms in Eq. 5 are part of the prior or likelihood, since it is only their product that matters. Thus, we can interpret stanhf models as having implicit flat priors for parameters between any bounds placed on them in the HistFactory model, and constraint terms that are part of the likelihood.
@@
Lastly, we may wish to choose a conjugate prior [45] — that is, a prior such that the prior and posterior belong to the same family of distributions. In this case, we may update using the auxiliary data as a preliminary step and specify only the posterior in the Stan model. That is, we could remove the constraint term and replace it with a posterior .
@@ III.6 Validation and testing
Checking that the parameter names and sizes agree; see stanhf.Converter.validate_par_names.
@@
The log-likelihoods are evaluated at randomly chosen points found by perturbing the suggested initial choices by noise. See stanhf.Converter.validate_target.
@@ IV Bayesian inference
The stanhf CLI builds an executable Stan program and writes associated data files. We briefly describe how to use them for Bayesian inference. For concreteness, we look here at the examples/normfactor.json model shown in Fig. 1. In this model, a signal is characterized by a strength parameter and corresponds to the background-only model. First, we must build the model and sample from it,
@@
We can plot posterior distributions. For example, to show the one-dimensional posterior pdf
[…]
Lastly, the Bayes factor [54] and Bayes factor surface [55] can be computed using a Savage-Dickey ratio [56]. The Bayes factor surface as a function of relative to the background-only model could be found by,
[…]
@@ V Frequentist inference
As well as using stanhf for Bayesian inference using the powerful HMC functionality of Stan, we can harness the expressive power of Stan models in a frequentist setting. stanhf provides classes and functions to use a Stan model and optimizer in place of a pyhf model or native pyhf optimizer in the statistical inference package of pyhf.infer, which uses results in ref. [57].333Frequentist inference is possible in Stan itself using simulation; see e.g. ref. [58].
As an example, we again use examples/normfactor.json, which declares that mu is the parameter of interest (POI). Using pyhf, we can plot the statistic [59] as a function of :
[…]
@@
VI Tweaking a model
Although the HistFactory declarative specification describes many common models, we may wish to go beyond it. For example, suppose we wish to model contributions to a signal, and , that depend on different powers of a modifier, e.g.,
This form could occur because of interference effects or because the contributions occur at different orders in perturbation theory.
We achieve this as follows. We begin from a model in which two signal contributions share a correlated normalization factor. This model can be found in the stanhf source code in examples/tweak.json. We generate a Stan model using stanhf examples/tweak.json. Because Stan is an imperative language, we can tweak the Stan code generated by stanhf as we wish. Thus, we edit the resulting tweak.stan model to match the required form Eq. 7. Specifically, we change the transformed parameters block such that the second contribution to the signal depends on the square of the normalization factor,
We show this process in full in Fig. 3.
VII Conclusions
We introduced stanhf — a new way to analyze and develop statistical models based on histogrammed data. The code transpiles HistFactory models into Stan, a probabilistic programming language. This allows state-of-the-art, scalable optimization and sampling using automatic differentiation and HMC. Because Stan is an imperative language, HistFactory modeling choices can be tweaked and adapted, as desired.
Stan is part of an ecosystem of Bayesian analysis software, allowing MCMC chains to be scrutinized for convergence and plotted, and even allowing models to be reran and analyzed online. Lastly, to ensure that the frequentist toolbox remains available, we provide an interface between Stan models and the inference machinery of pyhf.
@@ Acknowledgements.
- [3] O.A. Martin, R. Kumar and J. Lao, Bayesian Modeling and Computation in Python, Chapman and Hall/CRC (2021), DOI.
- [4] J. Brehmer, Simulation-based inference in particle physics, Nat. Rev. Phys. 3 (2021) 305 [2010.06439].
- [5] D.J. Lunn, A. Thomas, N. Best and D. Spiegelhalter, WinBUGS — A Bayesian modelling framework: Concepts, structure, and extensibility, Stat. Comput. 10 (2000) 325–337.
- [6] T.E. Fjelde, K. Xu, D. Widmann, M. Tarek, C. Pfiffer, M. Trapp et al., Turing.jl: a general-purpose probabilistic programming language, ACM Trans. Probab. Mach. Learn. (2025) .
- [8] D. Tran, A. Kucukelbir, A.B. Dieng, M. Rudolph, D. Liang and D.M. Blei, Edward: A library for probabilistic modeling, inference, and criticism, [1610.09787].
- [9] D. Tran, M.D. Hoffman, D. Moore, C. Suter, S. Vasudevan, A. Radul et al., Simple, Distributed, and Accelerated Probabilistic Programming in NeurIPS, 2018, URL.
- [10] E. Bingham, J.P. Chen, M. Jankowiak, F. Obermeyer, N. Pradhan, T. Karaletsos et al., Pyro: Deep Universal Probabilistic Programming, J. Mach. Learn. Res. 20 (2018) 1 [1810.09538].
- [11] D. Phan, N. Pradhan and M. Jankowiak, Composable Effects for Flexible and Accelerated Probabilistic Programming in NumPyro, [1912.11554].
@@
- [15] K. Cranmer, M. Drnevich, S. Macaluso and D. Pappadopulo, Reframing Jet Physics with New Computational Methods, EPJ Web Conf. 251 (2021) 03059 [2105.10512].
- [16] A. Güneş Baydin, L. Shao, W. Bhimji, L. Heinrich, L. Meadows, J. Liu et al., Etalumis: Bringing Probabilistic Programming to Scientific Simulators at Scale in International Conference for High Performance Computing, Networking, Storage, and Analysis (SC19), July, 2019 [1907.03382].
- [17] J. Qin and C. Tunnell, Approximate Differentiable Likelihoods for Astroparticle Physics Experiments in ACAT, 2024 [2408.09057].
- [18] A.G. Baydin, L. Shao, W. Bhimji, L. Heinrich, S. Naderiparizi, A. Munk et al., Efficient Probabilistic Inference in the Quest for Physics Beyond the Standard Model in NeurIPS, 2019, URL.
- [19] V.D. Elvira, S. Gottlieb, O. Gutsche, B. Nachman, S. Bailey, W. Bhimji et al., The Future of High Energy Physics Software and Computing in Computational Frontier Report Contribution to Snowmass 2021 [2210.05822].
@@
- [39] R. Češnovar, S. Bronder, D. Sluga, J. Demšar, T. Ciglarič, S. Talts et al., GPU-based Parallel Computation Support for Stan, [1907.01063].
- [40] J. Bradbury, R. Frostig, P. Hawkins, M.J. Johnson, C. Leary, D. Maclaurin et al., JAX: composable transformations of Python+NumPy programs, URL.
- [41] S. Maskell, Running Multiple Short MCMC Chains on a GPU Using JAX for Fast Inference with Stan in StanCon 2024 Oxford, URL.
@@
- [45] H. Raiffa and R. Schlaifer, Applied Statistical Decision Theory, John Wiley & Sons (1961).
@@
- [58] Stan Development Team, The Bootstrap and Bagging, URL.
- [59] A.L. Read, Presentation of search results: the technique, J. Phys. G: Nucl. Part. Phys. 28 (2002) 2693–2704.